The function f(x,y)={x^2y/x^4+y^2 (x,y) is not=0 0, (x,y)=0 is not continuous at (0,0)

1
Expert's answer
2022-05-04T13:20:06-0400

If "x=0" then "f(0, y)=\dfrac{0(y)}{0+y^2}=0." Therefore


"f(x, y)\to 0\text{ as} \ (x,y)\to(0,0) \text{ along the }y-\text{axis}"

For all "x\not=0"


"f(x, x^2)=\dfrac{x^2(x^2)}{x^4+(x^2)^2}=\dfrac{1}{2}\not=0"

"f(x, y)\to \dfrac{1}{2}\not=0\text{ as} \ (x,y)\to(0,0) \text{ along }y=x^2"

Since we have obtained different limits along different paths, limit


"\lim\limits_{(x,y)\to(0,0)}f(x, y)=\lim\limits_{(x,y)\to(0,0)}\dfrac{x^2y}{x^4+y^2}"

does not exist.

The function "f(x, y)=\dfrac{x^2y}{x^4+y^2}" is not continuous at "(0,0)."